# The cosmological population of Gamma-ray Bursts from the disks of Active Galactic Nuclei

### v1

@@ Abstract
With the discovery of gravitational waves (GWs), the disks of Active Galactic Nuclei (AGN) have emerged as an interesting environment for hosting a fraction of their sources. AGN disks are conducive to forming both long and short Gamma-Ray Bursts (GRBs), and their anticipated cosmological occurrence within these disks has potential to serve as an independent tool for probing and calibrating the population of stars and compact objects within them, and their contribution to the GW-detected population. In this study, we employ Monte Carlo methods in conjunction with models for GRB electromagnetic emission in extremely dense media to simulate the cosmological occurrence of both long and short GRBs within AGN disks, while also estimating their detectability across a range of wavelengths, from gamma-rays to radio frequencies. We investigate two extreme scenarios: “undiffused”, in which the radiation escapes without significant scattering (i.e. if the progenitor has excavated a funnel within the disk), and “diffused”, in which the radiation is propagated through the high-density medium, potentially scattered and absorbed. In the diffused case, we find that the majority of detectable GRBs are likely to originate from relatively low redshifts, and from the outermost regions of large supermassive black hole (SMBH) masses, $\gtrsim 10^{7.5}\rm M_{\odot}$. In the undiffused case, we expect a similar trend, but with a considerable contribution from the intermediate regions of lower SMBH masses. Detectable emission is generally expected to be dominant in prompt $\gamma$-rays if diffusion is not dominant, and X-ray afterglow if diffusion is important; however, the nature of the dominant observable signal highly depends on the specific AGN disk model, hence making GRBs in AGN disks also potential probes of the structure of the disks themselves.
@@ 1 Introduction
The study of Gamma-Ray Bursts (GRBs), with their long and short classes, has a history spanning more than three decades. Traditionally, these sources have been linked to galactic settings characterized by relatively modest ambient densities, $n$, ranging from a few to hundreds of atoms per cm3.
@@
The fact that both Long GRBs (LGRBs) and Short GRBs (SGRBs) could occur in AGN disks is not surprising. Stars are believed to exist within AGN disks, either forming in situ due to gravitational instabilities in the outer disk (e.g. Goodman 2003; Chen et al. 2023) or due to capture from the nuclear star cluster surrounding the AGN (e.g. Artymowicz et al. 1993; Fabj et al. 2020; Wang et al. 2024). The evolution of stars in these dense environments has been extensively studied recently (Cantiello et al., 2021; Jermyn et al., 2021; Dittmann et al., 2023; Chen et al., 2024; Liu et al., 2024). These studies have revealed that AGN stars, prior to going supernova, tend to grow to substantial masses and are characterized by high rotational speeds (but see (Chen & Lin, 2023), if strong disk turbulence is taken into account), making them ideal candidates as progenitors of LGRBs. Additionally, AGN disks are conducive to the formation of SGRBs, arising from binary mergers of neutron stars (NS) and NS-BH mergers with suitably small mass ratios. This propensity for compact object binary formation in AGN disks is due to the presence of migration traps, facilitating binary formation through dynamical interactions (Bellovary et al., 2016; Tagawa et al., 2020; Perna et al., 2021b; Grishin et al., 2024). Furthermore, the interaction of these objects with the dense AGN disk medium leads to the loss of kinetic energy, further promoting binary formation(Rowan et al., 2023; Li et al., 2023).
Recently a candidate for a GRB in a disk has been proposed (Levan et al. 2023; Lazzati et al. 2023, but see Stratta et al. 2024), lending credibility to these theoretical predictions. However, there has been no comprehensive analysis connecting the underlying AGN stellar population to the observable occurrence of both LGRBs and SGRBs. This paper aims to address this gap. Specifically, we employ Monte Carlo statistical techniques, coupled with models for the GRB electromagnetic emissions in highly dense media, to predict the properties of the population of observable GRBs originating from AGN disks. By providing theoretical predictions of prompt and afterglow emissions which can be compared with observations, we hope to: (a) Impose constraints on the stellar and compact object populations within AGN disks, and thus in turn help calibrate the proportion of binary BH merger events attributed to the AGN channel; (b) Quantify the contribution of GRB sources to the unknown variability observed in AGNs; (c) Identify features which are sensitive to the disk structure, and which can in turn help probe the structure of the AGN disk itself.
Figure 1: Redshift evolution of the global SMBH mass function. The mass profile has been reconstructed by interpolating the mass functions from Merloni & Heinz (2008).
@@ 2.1.2 Disk model
The Shakura-Sunyaev model (Shakura & Sunyaev 1973, SS in the following) has been remarkably successful in describing properties of accretion disks. However, it is known to have some shortcomings, such as being susceptible to gravitational instability in the outer regions. This has led to the development of increasingly more sophisticated models appropriate to the GN conditions, such as the one by Sirko & Goodman (2003) (SG from now on) and the one by Thompson et al. (2005) (TQM from now on), which enforce gravitational stability in the outer disk, and more recently by Gilbaum & Stone (2022) and Epstein-Martin et al. (2024), which include heating feedback in a more self-consistent way.
@@
The TQM model has six input parameters: SMBH mass, stellar dispersion, star formation efficiency, angular momentum efficiency, supernovae radiative fraction, and accretion rate. The stellar dispersion, $\sigma$, is calculated using the $M-\sigma$ relation. We adopt the standard TQM model values for star formation efficiency, $\epsilon_{T}=0.001$, angular momentum efficiency, $m=0.2$, and supernovae radiative fraction, $\xi=1$. Similar to SG, the accretion rate is not independently specified. However, unlike SG, the accretion rate in TQM is computed dynamically based on the disk’s physical properties and processes. The calculation begins with an initial accretion rate at the outer disk boundary, which evolves as the model integrates inward, taking into account radial mass transport and the effects of angular momentum transport. The accretion rate at a given radius is determined by: $\dot{M}=4\pi r\Omega_{T}m_{T}\rho h^{2}$ where $\rho$ is the gas density, $h$ is the disk scale height, $\Omega_{T}$ is the rotational velocity, and $m_{T}$ is the angular momentum efficiency. The Toomre parameter, $Q$, is used to evaluate disk stability; regions where $Q\sim 1$ undergo star formation, reducing $\dot{M}$ as mass is converted into stars. The star formation rate ($\dot{\Sigma}_{*}$) per unit area is given by: $\dot{\Sigma}_{*}=\Sigma_{g}\Omega_{T}\epsilon_{T}$ where $\Sigma_{g}$ is the surface mass density. Through this iterative process, the TQM model dynamically adjusts the accretion rate to account for star formation and feedback effects.
@@
Figure 2 illustrates disk properties of the SG and TQM models for a range of SMBH masses within $10^{6}$ $10^{9}$ M. The left plot shows the disk mid-plane number density, $n=\rho/m_{p}$ (where $m_{p}$ is the proton mass), as a function of disk radius in units of $R_{\rm g}$, with the gravitational radius defined as $R_{\rm g}=GM_{\rm SMBH}/c^{2}$, where $G$ is the gravity constant and $c$ is the speed of light. For both models, the number density is generally higher in the inner regions and at lower masses, while decreasing at large radii and for higher mass central BHs. However, the TQM model has a high-density localized region at distances within $10^{3}$ $10^{4}$ $\rm R_{g}$. This feature is closer to the central BH in the highest central-mass cases. The number density spans a wide range for both models, with the SG disk displaying a larger range than the TQM due to the steeper scaling in the outer regions. The right panels of Figure 2 show instead the scale height of the disk, $H$, which can approach $H/R\sim 1$ in the outermost regions, especially at high SMBH masses. A narrow, low scale-height feature is seen in the TQM model, mirroring the high density one discussed above.
@@ 2.2.1 Prompt emission
Figure 3 shows two sets of simulated light curves, one set for a short GRB engine (left panels) and another for a long one (right panels), illustrating both the dependence on the engine duration at fixed density (top panels) and on the density for fixed engine duration (bottom panels). For the chosen parameters (see caption of the figure), the transition from the standard, ’low-density’ GRBs to the HD regime generally happens between densities of $\sim 10^{5}-10^{8}$ cm-3, but with a precise value which depends on the engine duration, since we are assuming a fixed luminosity for all the bursts. For both the long and short GRB engines, above the critical density for transition, the light curves display a characteristic fast-rise slow-decay shape similar to the so-called FRED (fast rise, exponential decay) bursts (Fenimore et al., 1996). In addition, some overlaid variability is apparent in the brightest phases. Another characteristics found by Lazzati et al. (2022) is a monotonic hard-to-soft spectral evolution (Band et al., 1993), even during phases in which the burst is brightening. Of note, finally, is the fact that the short engine produces a burst that, given its duration, would be classified as long.
@@
A different look at the effect of high medium density on the prompt emission duration is taken in Figure 5. The probability distribution of the intrinsic engine duration, $T_{\rm eng}$, for both short and long GRBs is taken from observations (Bhat et al., 2016). Here we are making the reasonable assumptions that the physics of the central engine is not significantly dependent on the density at the burst location, and that the AGN-GRB sources are subdominant with respect to the global observed GRB population. Hence the properties of this population can be assumed to be a good proxy for the intrinsic properties of the GRB population in AGN disks. We fit the Bhat et al. (2016) data using a log-normal distribution to yield the number of bursts, $dn/dt_{\rm{eng}}$:
@@ 2.2.2 Afterglow emission
These frequencies have a different dependence on the density $n$ of the medium in which the fireball propagates: $\nu_{\rm inj}$ is independent of it, $\nu_{\rm c}$ is inversely proportional to the ambient density, while $\nu_{\rm abs}$ grows with it (e.g. Panaitescu & Mészáros 2000). This leads to important consequences: While in typical interstellar media densities the synchrotron frequency is smaller than the other two break frequencies, in very high density media it can grow to be the largest among the three, even reaching the X-ray band. As a result, the radio and optical/NIR afterglow fluxes are heavily suppressed compared to the standard afterglow in the ISM.
Examples of afterglow lightcurves in very dense media can be seen in Fig. 6 for four representative bands: Radio ($10^{9}$ Hz), Optical ($5\times 10^{14}$ Hz), X-rays ($10^{17}$ Hz), and Gamma rays ($10^{20}$ Hz). The top panels illustrate the dependence on the medium density at fixed viewing angle, for a sightly off-axis GRB, while the bottom panels show the dependence on the viewing angle for a fixed density. In all the cases, the GRB is a representative of a long GRB with isotropic energy $E_{\rm iso}=10^{53}$ ergs and initial Lorentz factor $\Gamma=300$. Notable is the high suppression of the radio flux with density, as a result of the $\nu^{2}$ low-frequency slope due to self-absorption. On the other hand, the synchrotron radiation produced by these electrons in gamma rays is above the self-absorption frequency, so the gamma ray flux density is not substantially suppressed (Granot & Sari, 2002; Piran, 2004). Note also that the emission peaks at earlier times for higher densities due to the fact the the jet break time occurs at earlier times for larger $n$, $t_{\rm peak}\propto n^{-1/3}$ (Sari et al., 1999).
@@ 3.1 Methods
We build the cosmological distribution of AGN-born GRBs using the model ingredients described in the previous section, with additional elements detailed below. Since this is the first of such works to date (to the best of our knowledge), we keep our model assumptions to be the simplest.
@@
(iii) Given the AGN disk with properties from (ii), we randomly draw the radial position $R_{\rm GRB}$ of the GRB within the AGN disk. Here, for both the case of LGRBS and SGRBs, we adopt a model-independent approach, in which the density of both LGRBs and SGRBs is proportional to the amount of mass at that location: $P(R)\propto P(\Sigma)$. By doing so, we assume GRBs are more likely to occur where the mass of the disk is densest, and that the distribution of GRBs is purely dependent on disk mass. This approach simplifies the physics to the most basic correlation between disk mass and GRB probability, making it a zeroth-order approximation of the true underlying processes. In future studies, we expect to use this as a baseline for comparison with more sophisticated models that involve more complex physical processes such as the effects of gravitational instabilities on triggering star formation, migration of stars, and feedback mechanisms from stellar physics. In particular, specific modeling of massive stars and of binary NS mergers in AGN disks suggests that migration traps or other dynamical processes may lead to preferential locations for GRB (Perna et al., 2021b; Fabj et al., 2024). These mechanisms, as well as their dependence on AGN disk properties, are still being investigated in various studies of AGN disks, and therefore we defer to follow up work incorporating them within our formalism.
@@
The last remaining variable of our simulated GRB population is the viewing angle $\theta_{\rm obs}$ of the jet axis with respect to the line of sight to the observer. We carry out prompt and afterglow radiation transfer calculations fixing this angle at $\theta_{\rm obs}=0$ to represent a face-on observation. Perna et al. (2021a) investigated the role of the viewing geometry on the location of the disk photosphere. They found that the viewing geometry plays a significant role only at large angles, when the line of sight intersects the outer parts of the disc ($\theta_{\rm{view}}\gtrsim 50^{o}$). At such large angles the disk photosphere increases dramatically, making the transients undetectable. Since our GRBs are assumed to be beamed within a narrow jet angle, however, bursts at large viewing angles would be intrinsically undetectable even in the undiffused case.

### v2

@@ Abstract
With the discovery of gravitational waves (GWs), the disks of Active Galactic Nuclei (AGN) have emerged as an interesting environment for hosting a fraction of their sources. AGN disks are conducive to forming both long and short Gamma-Ray Bursts (GRBs), and their anticipated cosmological occurrence within these disks has potential to serve as an independent tool for probing and calibrating the population of stars and compact objects within them, and their contribution to the GW-detected population. In this study, we employ Monte Carlo methods in conjunction with models for GRB electromagnetic emission in extremely dense media to simulate the cosmological occurrence of both long and short GRBs within AGN disks, while also estimating their detectability across a range of wavelengths, from gamma-rays to radio frequencies. We investigate two extreme scenarios: “undiffused”, in which the radiation escapes without significant scattering (i.e. if the progenitor has excavated a funnel within the disk), and “diffused”, in which the radiation is propagated through the high-density medium, potentially scattered and absorbed. In the diffused case, we find that the majority of detectable GRBs, which are at most a few percent of the total, are likely to originate from relatively low redshifts, and from the outermost regions of large supermassive black hole (SMBH) masses, $\gtrsim 10^{7.5}\rm M_{\odot}$. In the undiffused case, which has a GRB detection probability $\sim 40-50\%$, we expect a similar trend, but with a considerable contribution from the intermediate regions of lower SMBH masses. Detectable emission is generally expected to be dominant in prompt $\gamma$-rays if diffusion is not dominant, and X-ray afterglow if diffusion is important; however, the nature of the dominant observable signal highly depends on the specific AGN disk model, hence making GRBs in AGN disks also potential probes of the structure of the disks themselves.
@@ 1 Introduction
The study of Gamma-Ray Bursts (GRBs), with their long and short classes, has a history spanning more than three decades. Traditionally, these sources have been linked to galactic environments characterized by relatively modest ambient densities, $n$, ranging from a few to hundreds of atoms per cm3 (Scalo & Wheeler, 2001; Holland et al., 2002; Muccino et al., 2013; Krongold & Prochaska, 2013; Fong et al., 2015).
@@
The fact that both Long GRBs (LGRBs) and Short GRBs (SGRBs) could occur in AGN disks is not surprising. Stars are believed to exist within AGN disks, either forming in situ due to gravitational instabilities in the outer disk (e.g. Goodman 2003; Chen et al. 2023) or due to capture from the nuclear star cluster surrounding the AGN (e.g. Artymowicz et al. 1993; Fabj et al. 2020; Wang et al. 2024). The evolution of stars in these dense environments has been extensively studied recently (Cantiello et al., 2021; Jermyn et al., 2021; Dittmann et al., 2023; Chen et al., 2024; Liu et al., 2024). These studies have revealed that AGN stars, prior to going supernova, tend to grow to substantial masses and are characterized by high rotational speeds (but see Chen & Lin 2023, if strong disk turbulence is taken into account), making them ideal candidates as progenitors of LGRBs. Additionally, AGN disks are conducive to the formation of SGRBs, arising from binary mergers of neutron stars (NS) and NS-BH mergers with suitably small mass ratios. This propensity for compact object binary formation in AGN disks is due to the presence of migration traps, facilitating binary formation through dynamical interactions (Bellovary et al., 2016; Tagawa et al., 2020; Perna et al., 2021b; Grishin et al., 2024). Furthermore, the interaction of these objects with the dense AGN disk medium leads to the loss of kinetic energy, further promoting binary formation(Rowan et al., 2023; Li et al., 2023).
Recently a candidate for a GRB in a disk has been proposed (Levan et al. 2023; Lazzati et al. 2023, but see Stratta et al. 2024), lending credibility to these theoretical predictions [^1] Note also the case of GRB 150101B hosted in a galaxy with an AGN (Xie et al., 2016). However, its offset from the centre does not make this GRB a candidate for an AGN-disk host.. However, there has been no comprehensive analysis connecting the underlying AGN stellar population to the observable occurrence of both LGRBs and SGRBs. This paper aims to address this gap. Specifically, we employ Monte Carlo statistical techniques, coupled with models for the GRB electromagnetic emissions in highly dense media, to predict the properties of the population of observable GRBs originating from AGN disks. Even though this population is expected to be largely subdominant with respect to the global observed GRB population, by providing theoretical predictions of its prompt and afterglow emission which can be compared with observations, we hope to: (a) Impose constraints on the stellar and compact object populations within AGN disks, and thus in turn help calibrate the proportion of binary BH merger events attributed to the AGN channel; (b) Quantify the contribution of GRB sources to the unknown variability observed in AGNs; (c) Identify features which are sensitive to the disk structure, and which can in turn help probe the structure of the AGN disk itself.
Figure 1: Redshift evolution of the global SMBH mass function density. The mass profile has been reconstructed by interpolating the mass functions from Merloni & Heinz (2008).
@@ 2.1.2 Disk model
The Shakura-Sunyaev model (Shakura & Sunyaev 1973, SS in the following) has been remarkably successful in describing properties of accretion disks. However, it is known to have some shortcomings, such as being susceptible to gravitational instability in the outer regions. This has led to the development of increasingly more sophisticated models appropriate to the AGN conditions, such as the one by Sirko & Goodman (2003) (SG from now on) and the one by Thompson et al. (2005) (TQM from now on), which enforce gravitational stability in the outer disk, and more recently by Gilbaum & Stone (2022) and Epstein-Martin et al. (2024), which include heating feedback in a more self-consistent way.
@@
The TQM model has six input parameters: SMBH mass, stellar dispersion, star formation efficiency, angular momentum efficiency, supernovae radiative fraction, and accretion rate. The stellar dispersion, $\sigma$, is calculated using the $M-\sigma$ relation. We adopt the standard TQM model values for star formation efficiency, $\epsilon_{T}=0.001$, angular momentum efficiency, $m=0.2$, and supernovae radiative fraction, $\xi=1$. Similar to SG, the accretion rate is not independently specified. However, unlike SG, the accretion rate in TQM is computed dynamically based on the disk’s physical properties and processes. The calculation begins with an initial accretion rate at the outer disk boundary, which evolves as the model integrates inward, taking into account radial mass transport and the effects of angular momentum transport. The accretion rate at a given radius is determined by: $\dot{M}=4\pi r\Omega_{T}m_{T}\rho H^{2}$ where $\rho$ is the gas density, $H$ is the disk scale height, $\Omega_{T}$ is the rotational velocity, and $m_{T}$ is the angular momentum efficiency (Thompson et al., 2005). The Toomre parameter, $Q$, is used to evaluate disk stability; regions where $Q\sim 1$ undergo star formation, reducing $\dot{M}$ as mass is converted into stars. The star formation rate ($\dot{\Sigma}_{*}$) per unit area is given by: $\dot{\Sigma}_{*}=\Sigma_{g}\Omega_{T}\epsilon_{T}$ where $\Sigma_{g}$ is the surface mass density. Through this iterative process, the TQM model dynamically adjusts the accretion rate to account for star formation and feedback effects.
@@
Figure 2 illustrates disk properties of the SG and TQM models for a range of SMBH masses within $10^{6}$ $10^{9}$ M. The left plot shows the disk mid-plane number density, $n=\rho/m_{p}$ (where $m_{p}$ is the proton mass), as a function of both disk radius (in units of $R_{\rm g}$) and SMBH mass (in units of M). The gravitational radius is defined as $R_{\rm g}=GM_{\rm SMBH}/c^{2}$, where $G$ is the gravity constant and $c$ is the speed of light. For both models, the number density is generally higher in the inner regions and at lower masses, while decreasing at large radii and for higher mass central BHs. The sharp peak around $10^{3}$ $10^{4}$ $R_{\rm g}$ in the TQM model is due to dominant gravitational instability as the Toomre parameter, $Q$, approaches 1. This leads to rapid cooling, allowing gas to accumulate. Beyond $10^{4}$ $R_{\rm g}$, star formation feedback regulates the disk, causing the density to decline. On the other hand, the SG model gradually approaches gravitational instability, resulting in two moderate density peaks. The inner peak corresponds to the transition from the radiation pressure-dominated regime to the gas pressure-dominated regime. The second peak around $10^{3}$ $R_{\rm g}$ is the region where self-gravity starts becoming significant, leading to enhanced mass accumulation, but the disk remains only marginally gravitationally unstable. The number density spans a wide range for both models, with the SG disk displaying a larger range than the TQM due to the steeper scaling in the outer regions. The right panels of Figure 2 show instead the scale height of the disk, $H$, which can approach $H/R\sim 1$ in the outermost regions, especially at high SMBH masses. A narrow, low scale-height feature is seen in the TQM model, mirroring the high density one discussed above.
@@ 2.2.1 Prompt emission
Figure 3 shows two sets of simulated light curves, one set for a short GRB engine (left panels) and another for a long one (right panels), illustrating both the dependence on the engine duration at fixed density (top panels) and on the density for fixed engine duration (bottom panels). In all cases the code described in Lazzati et al. (2022) was used. The code assumes that the outflow is made of a fixed number of shells (five, in our case) with low Lorentz factor dispersion ($\Delta\Gamma/\Gamma\sim 2$). If the external density is above the value specified in Eq 2, the shell dynamics is computed taking into account that the first shell interacts with the static, uniform external medium, while subsequent shells interact with the system made by the decelerated previous shells plus the shocked ambient material. The efficiency is calculated self consistently by conserving energy and momentum and the duration of each pulse is taken to be dominated by the curvature effect. If, instead, the density is below the critical value, the emission is due to collisions between pairs of shells before any interaction with the ambient medium. Collision radii are calculated with basic kinematics and the efficiency is derived from energy and momentum conservation, like in the previous case. Pulse duration is, similarly, dominated by the curvature effect. Each pulse is given a double exponential shape, with the decay phase lasting longer than the rising one. For the chosen parameters (see caption of the figure), the transition from the standard, ’low-density’ GRBs to the HD regime generally happens between densities of $\sim 10^{5}-10^{8}$ cm-3, but with a precise value which depends on the engine duration, since we are assuming a fixed luminosity for all the bursts. For both the long and short GRB engines, above the critical density for transition, the light curves display a characteristic fast-rise slow-decay shape similar to the so-called FRED (fast rise, exponential decay) bursts (Fenimore et al., 1996). In addition, some overlaid variability is apparent in the brightest phases. Another characteristics found by Lazzati et al. (2022) is a monotonic hard-to-soft spectral evolution (Band et al., 1993), even during phases in which the burst is brightening. Of note, finally, is the fact that the short engine produces a burst that, given its duration, would be classified as long.
@@
A different look at the effect of high medium density on the prompt emission duration is taken in Figure 5. The probability distribution of the intrinsic engine duration, $T_{\rm eng}$, for both short and long GRBs is assumed to be the same as the $T_{90}$ distribution [^2] While this equality is not strictly true, it is a necessary approximation here given that the exact relation varies from burst to burst depending on the spectrum and the observational band and instrumental sensitivity., with the latter taken from observations from the 3rd Fermi/GBM GRB catalogue (Bhat et al., 2016). Here we are making the reasonable assumptions that the physics of the central engine is not significantly dependent on the density at the burst location, and that the AGN-GRB sources are subdominant with respect to the global observed GRB population. Hence the properties of this population can be assumed to be a good proxy for the intrinsic properties of the GRB population in AGN disks. We fit the Bhat et al. (2016) data using a log-normal distribution to yield the number of bursts, $dn/dt_{\rm{eng}}$:
@@ 2.2.2 Afterglow emission
These frequencies have a different dependence on the density $n$ of the medium in which the fireball propagates: $\nu_{\rm inj}$ is independent of it, $\nu_{\rm c}$ is inversely proportional to the ambient density, while $\nu_{\rm abs}$ grows with it (e.g. Panaitescu & Mészáros 2000). This leads to important consequences: While in typical interstellar media densities the synchrotron self-absorption frequency $\nu_{\rm abs}$ is smaller than the other two break frequencies, in very high density media it can grow to be the largest among the three, even reaching the X-ray band. As a result, the radio and optical/NIR afterglow fluxes are heavily suppressed compared to the standard afterglow in the ISM. Since the code is designed to provide input for diffusion calculations (Wang et al., 2022) the light curves are generated using a Monte Carlo method. Photons are seeded in the external shock according to the local emissivity, isotropic in the comoving frame. Their direction and frequency is then boosted in the laboratory frame. Light curves are then generated by collecting and binning the arrival times of photons in the direction of the selected observer.
Examples of afterglow lightcurves in very dense media can be seen in Fig. 6 for four representative bands: Radio ($10^{9}$ Hz), Optical ($5\times 10^{14}$ Hz), X-rays ($10^{17}$ Hz), and Gamma rays ($10^{20}$ Hz). The top panels illustrate the dependence on the medium density at fixed viewing angle, for a sightly off-axis GRB, while the bottom panels show the dependence on the viewing angle for a fixed density. In all the cases, the GRB is a representative of a long GRB with isotropic-equivalent energy $E_{\rm iso}=10^{53}$ ergs and initial Lorentz factor $\Gamma=300$. The jet geometry is assumed to be top-hat, with an opening angle of $5^{\circ}$. Ten million photons were used to generate each set of light curves. Still, the effect of a finite photon number causes the noisy behavior at very low fluxes, easily seen in the upper center and right panels. Notable is the high suppression of the radio flux with density, as a result of the $\nu^{2}$ low-frequency slope due to self-absorption. On the other hand, the synchrotron radiation produced by these electrons in gamma rays is above the self-absorption frequency, so the gamma ray flux density is not substantially suppressed (Granot & Sari, 2002; Piran, 2004). Note also that the emission peaks at earlier times for higher densities due to the fact the jet break time occurs at earlier times for larger $n$, $t_{\rm peak}\propto n^{-1/3}$ (Sari et al., 1999).
@@ 3.1 Methods
We build the cosmological distribution of AGN-born GRBs using the model ingredients described in the previous sections, with additional elements detailed below. Since this is the first of such works to date (to the best of our knowledge), we keep our model assumptions to be the simplest.
@@
(iii) Given the AGN disk with properties from (ii), we randomly draw the radial position $R_{\rm GRB}$ of the GRB within the AGN disk. Here, for both the case of LGRBS and SGRBs, we adopt a model-independent approach, in which the density of both LGRBs and SGRBs is proportional to the surface mass density at that radius: $P(R)\propto P(\Sigma)$. By doing so, we assume GRBs are more likely to occur where the surface mass density of the disk is highest, and that the distribution of GRBs is purely dependent on the local surface mass density where disk mass is most concentrated. This approach simplifies the physics to the most basic correlation between disk mass and GRB probability, making it a zeroth-order approximation of the true underlying processes. In future studies, we expect to use this as a baseline for comparison with more sophisticated models that involve more complex physical processes such as the effects of gravitational instabilities on triggering star formation, migration of stars, and feedback mechanisms from stellar physics. In particular, specific modeling of massive stars and of binary NS mergers in AGN disks suggests that migration traps or other dynamical processes may lead to preferential locations for GRB (Perna et al., 2021b; Fabj et al., 2024). These mechanisms, as well as their dependence on AGN disk properties, are still being investigated in various studies of AGN disks, and therefore we defer to follow up work incorporating them within our formalism.
@@
The last remaining variable of our simulated GRB population is the viewing angle $\theta_{\rm obs}$ of the jet axis with respect to the line of sight to the observer. We carry out prompt and afterglow radiation transfer calculations fixing this angle at $\theta_{\rm obs}=0$ to represent a face-on observation. Perna et al. (2021a) investigated the role of the viewing geometry on the location of the disk photosphere. They found that the viewing geometry plays a significant role only at large angles, when the line of sight intersects the outer parts of the disc ($\theta_{\rm{obs}}\gtrsim 50^{o}$). At such large angles the disk photosphere increases dramatically, making the transients undetectable. Since our GRBs are assumed to be beamed within a narrow jet angle, however, bursts at large viewing angles would be intrinsically undetectable even in the undiffused case.